# The Invisible Leash? Why RLVR May or May Not Escape Its Origin

### v3

@@ 3.2 Results: Predominant Preservation with Limited Expansion
Table 1 shows that across diverse model scales and families, RLVR predominantly acts as a support-constrained optimization mechanism. All models achieve very high support retention (overall $\text{SRR}\approx 0.93$$0.99$) while genuine discovery remains rare ($\text{NDR}\leq 0.04$). For example, Nemotron-7B and Nemotron-14B retain nearly all base-model solutions ($\text{SRR}\geq 0.98$) with negligible expansion. Even smaller-scale ProRL-1.5B achieves $\text{SRR}=0.93$ with only modest gains ($\text{NDR}=0.02$). These patterns persist across math ($\text{SRR}=0.96$$0.99$, $\text{NDR}\approx 0.00$$0.01$) and non-math domains ($\text{SRR}=0.90$$0.99$, $\text{NDR}\leq 0.04$). Preservation-dominated behavior is especially clear in Reasoning Gym tasks such as graph_color and palindrome, where ProRL accelerates convergence toward near-perfect pass@k with large budgets (Fig. 1). Support counts confirm this: most correct completions remain shared between RLVR and base models.
@@
Despite the strong conservation, RLVR occasionally recovers solutions negligible to the base model. Expansion is consistently small: ProRL-1.5B discovers 48 new completions across 11 benchmarks, while larger models (e.g., Phi4-14B, Nemotron-14B) add fewer than 10. Non-math datasets exhibit the highest relative discovery ($\text{NDR}\leq 0.04$), whereas math datasets are virtually stagnant ($\text{NDR}\leq 0.01$). Some Reasoning Gym tasks, such as graph_color_vertex20 and arc_1d, show genuine expansion (Fig. 2), but remain isolated exceptions rather than the dominant trend. These suggest that while RLVR can occasionally redistribute mass into underexplored solution modes, such expansion remains the exception rather than the rule, challenging assumptions about RLVR’s capacity for genuine reasoning horizon extension.
@@
Across all models and domains, shrinkage consistently exceeds expansion. ProRL-1.5B loses 175 completions while gaining only 48 (ratio $\approx 3.6{:}1$), while Nemotron-7B and Skywork-OR1-7B display similar patterns (ratios $\approx 2{:}1$3:13{:}1). Even large models (Nemotron-14B, Phi4-14B) show net shrinkage despite near-perfect preservation. Overall NSCR values remain slightly negative (0.01-0.01 to 0.06-0.06), showing that RLVR systematically narrows the accessible solution set. This explains paradoxical outcomes: while RLVR models outperform bases at low $k$, base models dominate at high $k$ due to broader solution coverage (e.g., AIME2024 base $\texttt{pass@8192}=93.3\%$ vs. ProRL-1.5B’s $83.3\%$). Reasoning Gym tasks like leg_counting, family_relationships, and power_function illustrate this vividly (Fig. 3).
@@
Support dynamic score confirms imbalance. SDS values remain consistently low across all scales ($\leq 0.07$), reflecting poor balance between preservation and discovery. The highest observed SDS is LiveBench-L at $0.288$, but even this corresponds to 2 expansions against 4 shrinkages. Math benchmarks are particularly imbalanced (SDS $\approx 0.00$$0.01$), while non-math domains fare only marginally better. Thus, RLVR’s improvements arise primarily from mass concentration, not meaningful solution expansion.
@@
Perplexity analysis on support constraints. Tab. 3 reports perplexity where base and ProRL models are evaluated against external reasoning traces from DeepSeek-R1 and Claude Sonnet 4. ProRL consistently shows higher perplexity. For instance, on AIME2024, perplexity against Claude Sonnet 4 rises from 8.76 (Base) to 14.91 (ProRL), which indicates RLVR reduces the model’s ability to assign probability to diverse external reasoning styles. While format differences contribute, the dominant effect is structural: RLVR concentrates probability around narrower solution trajectories. Correctness patterns further reveal the precision-coverage trade-off. In shrinkage cases, ProRL shows higher perplexity even when the base model succeeds, confirming RLVR collapses mass away from viable solution pathways. Conversely, modest perplexity improvements in rare expansion cases suggest ProRL’s new successful trajectories originate from the base model’s low-density tails rather than genuinely novel reasoning structures.
Dataset Training pass@k Support Dynamics Metrics Support Counts
Base Target SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@256)
AIME2024 SFT 63.33% 63.33% 0.789 0.211 0.332 0.000 15 4 4 7
DAPO 63.33% 66.67% 0.895 0.150 0.257 0.045 17 3 2 8
AIME2025 SFT 50.00% 63.33% 0.933 0.263 0.411 0.200 14 5 1 10
DAPO 50.00% 56.67% 1.000 0.118 0.211 0.118 15 2 0 13
AMC23 SFT 100.00% 100.00% 1.000 0.000 0.000 0.000 40 0 0 0
DAPO 100.00% 92.50% 0.925 0.000 0.000 -0.075 37 0 3 0
MATH500 SFT 96.40% 99.00% 0.996 0.030 0.059 0.026 480 15 2 3
DAPO 96.40% 95.20% 0.975 0.013 0.025 -0.012 470 6 12 12
Minerva SFT 63.24% 66.91% 0.895 0.154 0.263 0.050 154 28 18 72
DAPO 63.24% 55.15% 0.802 0.080 0.145 -0.120 138 12 34 88
Olympiad SFT 78.22% 81.78% 0.953 0.089 0.162 0.042 503 49 25 98
DAPO 78.22% 72.89% 0.884 0.051 0.096 -0.065 467 25 61 122
@@
Comparison between SFT and RLVR. We fix the base model (Qwen2.5-Math-7B), dataset (DeepMath-103K), sampling protocol, and optimization hyperparameters, varying only the training objective (SFT vs. DAPO). As shown in Tab. 2 and 12, SFT consistently produces moderate support expansion with positive NSCR values across benchmarks, whereas DAPO produces sharply concentrated distributions with mixed or negative support change. While DAPO improves precision at low sampling budgets, its support dynamics reflect a high-SRR but low-NDR regime: preservation of known correct solutions but limited discovery of new ones. On several benchmarks such as MATH500, Minerva, and Olympiad, DAPO reduces accessible correct solutions, causing net shrinkage despite identical training conditions. These results confirm that support-constrained behavior is not an artifact of scaling, data mixture, or procedure, but emerges from the objective itself. By contrast, SFT expands the reachable solution set without collapsing existing modes, underscoring that shrinkage is intrinsic to RLVR-style multiplicative updates rather than SFt.
Support dynamics during RLVR training. Tab, 13 summarizes how support-dynamics metrics evolve across RLVR training steps. Across benchmarks, SRR remains consistently high, showing that RLVR reliably preserves previously correct reasoning paths, while NSCR gradually decreases, indicating a steady contraction of the model’s solution support. On harder datasets, mild fluctuations in NDR and SDS reflect transient exploration that is subsequently pruned. Overall, these checkpoint-level statistics reveal that RLVR reshapes rather than monotonically improves the reasoning distribution. It reinforces a narrow set of stable trajectories over time, which explains both the early-stage gains and the late-stage degradation observed on more diverse math tasks.
Overall takeaway: RLVR as precision enhancer, not capability expander. Across model scales (1.5B–14B) and domains (math, non-math, multimodal), RLVR consistently behaves as a support-bounded optimizer. With SRR near one but NDR near zero, and uniformly negative NSCR, RLVR enhances precision by concentrating mass on known high-reward solutions but rarely discovers new reasoning paths. This aligns with the Temporal Forgetting effect (Li et al., 2025). Breaking RLVR’s invisible leash may thus require explicit exploration or hybrid strategies that deliberately seed probability mass into underrepresented solution regions.
@@
Category Correctness Reference Target AIME 2024 AIME 2025 Olympiad
@@ 3.3 When Empirical Support Expansion Occurs
(1) RLVR Recomposes Subskills the Base Model Already Possesses
@@
(2) RLVR Corrects Prompt-Format Misalignment and Extracts Latent Ability
A second driver is the sensitivity to prompt format. In several expanding tasks, especially dice and boxnet, the base model demonstrates partial competence but fails to follow the exact response format required by the reward function. RLVR reshapes the distribution to follow instructions more effectively, unlocking capabilities that were previously present but suppressed. Consistent with this interpretation, the perplexity gaps in Table 3 remain modest in expansion cases, showing that the “new” completions are stylistic or formatting variants of reasoning patterns already accessible to the base model, not fundamentally novel solutions beyond its support.
Why Expansion Is Small and Bounded
Even when these two favorable mechanisms align, expansion remains sharply limited. Across all benchmarks, NDR never exceeds $0.04$ (Tab. 1), NSCR remains negative in all models. Thus, RLVR’s gains arise from amplifying low-probability but existing solution fragments or format-correct variants, but never from discovering solutions truly absent from the base distribution.
4 Entropy Reduction and the pass@k Trade-off
4.1 Experimental Setup
To study how RLVR reshapes the sampling distribution, we examine the base model and RLVR with a medium sampling budget $k=32$ on the math reasoning benchmarks. We quantify changes in the output distribution using two entropy metrics:
4.2 Results: Precision Gains, Entropy Dynamics, and Trade-offs
@@
Metric Model AIME 2024 AMC 2023 MATH 500 Minerva Olympiad Avg.
@@
Tab. 4 shows that RLVR consistently improves avg@32 across all benchmarks, raising average performance from 54.5% to 65.4% for ProRL and from 43.0% to 61.3% for DAPO (Yu et al., 2025). However, this increased precision comes at a cost: RLVR systematically reduces answer-level entropy, indicating a collapse onto fewer distinct solutions and empirically validating our theoretical prediction that reward optimization sharpens output distributions around known modes, thereby reducing effective support coverage. Notably, intrinsically harder tasks, such as AIME or Minerva, still exhibit higher absolute answer-level entropy for both the base and RLVR models, suggesting that challenging problems inherently foster broader solution spaces that require exploration over more diverse completions.
@@
Our empirical analysis reveals a trade-off in RLVR: it improves precision by amplifying high-reward outputs, but simultaneously narrows the diversity of global solutions. This limitation is especially consequential in domains that admit multiple valid answers or benefit from creative reasoning, underscoring the need for explicit exploration mechanisms or diversity-promoting strategies to complement standard RLVR. Moreover, the observed divergence between token-level and answer-level entropy highlights the need for a more nuanced interpretation of stochasticity in reward-optimized models—showing that precision gains often come at the expense of global diversity, and that maintaining controlled variability is critical for sustaining effective exploration.

### v4

@@ 4.2 Results: Predominant Preservation with Limited Expansion
Tab. 1 shows that across diverse model scales and families, RLVR predominantly acts as a support-constrained optimization mechanism. All 1.5B–14B models achieve very high support retention (overall $\text{SRR}\approx 0.93$$0.99$) while genuine discovery remains rare ($\text{NDR}\leq 0.04$). For example, Nemotron-7B and Nemotron-14B retain nearly all base-model solutions ($\text{SRR}\geq 0.98$) with negligible expansion. Even smaller-scale ProRL-1.5B-v2 achieves $\text{SRR}=0.93$ with only modest gains ($\text{NDR}=0.02$). These patterns persist across math ($\text{SRR}=0.96$$0.99$, $\text{NDR}\approx 0.00$$0.01$) and non-math domains ($\text{SRR}=0.90$$0.99$, $\text{NDR}\leq 0.04$). Preservation-dominated behavior is especially clear in Reasoning Gym tasks such as graph_color and palindrome, where ProRL accelerates convergence toward near-perfect pass@k with large budgets (Fig. 2). Support counts confirm this: most correct completions remain shared between RLVR and base models.
@@
Despite strong conservation, RLVR occasionally recovers solutions that are negligible relative to the base model. Expansion is consistently small: ProRL-1.5B-v2 discovers 48 new completions across 11 benchmarks, while larger models (e.g., Phi4-14B, Nemotron-14B) add fewer than 10 per domain. Non-math datasets exhibit the highest relative discovery ($\text{NDR}\leq 0.04$), whereas math datasets are virtually stagnant ($\text{NDR}\leq 0.01$). Some Reasoning Gym tasks, such as graph_color_vertex20 and arc_1d, show genuine expansion (Fig. 3), but remain isolated exceptions rather than the dominant trend. These suggest that while RLVR can occasionally redistribute mass into underexplored solution modes, such expansion remains the exception rather than the rule, challenging assumptions about RLVR’s capacity for genuine reasoning horizon extension.
@@
Across all models and domains, shrinkage consistently exceeds expansion. ProRL-1.5B-v2 loses 175 completions while gaining only 48 (ratio $\approx 3.6{:}1$), while Nemotron-7B and Skywork-OR1-7B display similar patterns (ratios $\approx 2{:}1$3:13{:}1). Even large models (Nemotron-14B, Phi4-14B) show net shrinkage despite near-perfect preservation. Overall NSCR values remain slightly negative (0.01-0.01 to 0.06-0.06), showing that RLVR systematically narrows the accessible solution set. This explains paradoxical outcomes: while RLVR models outperform bases at low $k$, base models dominate at high $k$ due to broader solution coverage (e.g., AIME2024 base $\texttt{pass@8192}=93.3\%$ vs. ProRL-1.5B’s $83.3\%$). Reasoning Gym tasks like leg_counting, family_relationships, and power_function illustrate this vividly (Fig. 4).
@@
Support dynamic score confirms imbalance. SDS values remain consistently low in the domain-level aggregates for our 1.5B–14B models ($\leq 0.07$), reflecting poor balance between preservation and discovery. TFor example, LiveBench-L reaches SDS = $0.278$ (2 expansions vs. 2 shrinkages). Math benchmarks are particularly imbalanced (SDS $\approx 0.00$$0.01$), while non-math domains fare only marginally better. Thus, RLVR’s improvements arise primarily from mass concentration rather than meaningful solution expansion.
@@
Perplexity analysis on support constraints. Tab. 2 reports perplexity where base and ProRL models are evaluated against external reasoning traces from DeepSeek-R1 and Claude Sonnet 4. ProRL consistently shows higher perplexity. For instance, on AIME2024, perplexity on Claude Sonnet 4 increases from 8.76 (Base) to 14.91 (ProRL), indicating that RLVR reduces the model’s ability to assign probabilities to diverse external reasoning styles. While format differences contribute, the dominant effect is structural: RLVR concentrates probability around narrower solution trajectories. Correctness patterns further reveal the precision-coverage trade-off. In shrinkage cases, ProRL shows higher perplexity even when the base model succeeds, confirming RLVR collapses mass away from viable solution pathways. Conversely, modest perplexity improvements in rare expansion cases suggest ProRL’s new successful trajectories originate from the base model’s low-density tails rather than genuinely novel reasoning structures.
@@
Comparison between SFT and RLVR. We fix the base model (Qwen2.5-Math-7B), dataset (DeepMath-103K), sampling protocol, and optimization hyperparameters, varying only the training objective (SFT vs. DAPO). As shown in Tab. 3 and 14, SFT consistently produces moderate support expansion with positive NSCR values across benchmarks, whereas DAPO produces sharply concentrated distributions with mixed or negative support change. While DAPO improves precision at low sampling budgets, its support dynamics reflect a high-SRR but low-NDR regime: it preserves known correct solutions but discovers few new ones. On several benchmarks, such as MATH500, Minerva, and Olympiad, DAPO reduces the number of accessible correct solutions, resulting in net shrinkage despite identical training conditions. These results confirm that support-constrained behavior is not an artifact of scaling, data mixture, or procedure, but emerges from the objective itself. By contrast, SFT expands the reachable solution set without collapsing existing modes, underscoring that shrinkage is intrinsic to RLVR-style multiplicative updates rather than SFT.
@@
Category Correctness Reference Target AIME24 AIME25 Olympiad
@@
Dataset Training pass@k Support Dynamics Metrics Support Counts
Base Target SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@256)
AIME24 SFT 63.33% 63.33% 0.789 0.211 0.332 0.000 15 4 4 7
DAPO 63.33% 66.67% 0.895 0.150 0.257 0.045 17 3 2 8
AIME25 SFT 50.00% 63.33% 0.933 0.263 0.411 0.200 14 5 1 10
DAPO 50.00% 56.67% 1.000 0.118 0.211 0.118 15 2 0 13
AMC23 SFT 100.00% 100.00% 1.000 0.000 0.000 0.000 40 0 0 0
DAPO 100.00% 92.50% 0.925 0.000 0.000 -0.075 37 0 3 0
MATH500 SFT 96.40% 99.00% 0.996 0.030 0.059 0.026 480 15 2 3
DAPO 96.40% 95.20% 0.975 0.013 0.025 -0.012 470 6 12 12
Minerva SFT 63.24% 66.91% 0.895 0.154 0.263 0.050 154 28 18 72
DAPO 63.24% 55.15% 0.802 0.080 0.145 -0.120 138 12 34 88
Olympiad SFT 78.22% 81.78% 0.953 0.089 0.162 0.042 503 49 25 98
DAPO 78.22% 72.89% 0.884 0.051 0.096 -0.065 467 25 61 122
Dataset Step Pass@256 Support Dynamics Metrics Support Counts Dataset Step Pass@256 Support Dynamics Metrics Support Counts
RLVR SRR NDR SDS NSCR P E S O RLVR SRR NDR SDS NSCR P E S O
AIME24 0 83.33 - - - - - - - - AIME25 0 70.00 - - - - - - - -
30 80.00 0.960 0.000 0.000 -0.040 24 0 1 5 30 70.00 0.905 0.095 0.172 0.000 19 2 2 7
60 80.00 0.960 0.000 0.000 -0.040 24 0 1 5 60 73.33 0.905 0.136 0.237 0.042 19 3 2 6
90 83.33 1.000 0.000 0.000 0.000 25 0 0 5 90 70.00 0.905 0.095 0.172 0.000 19 2 2 7
120 80.00 0.960 0.000 0.000 -0.040 24 0 1 5 120 73.33 0.905 0.136 0.237 0.042 19 3 2 6
150 80.00 0.960 0.000 0.000 -0.040 24 0 1 5 150 73.33 0.905 0.136 0.237 0.042 19 3 2 6
180 83.33 1.000 0.000 0.000 0.000 25 0 0 5 180 63.33 0.905 0.000 0.000 -0.095 19 0 2 9
210 76.67 0.920 0.000 0.000 -0.080 23 0 2 5 210 66.67 0.905 0.050 0.095 -0.045 19 1 2 8
240 73.33 0.880 0.000 0.000 -0.120 22 0 3 5 240 66.67 0.905 0.050 0.095 -0.045 19 1 2 8
270 73.33 0.880 0.000 0.000 -0.120 22 0 3 5 270 63.33 0.905 0.000 0.000 -0.095 19 0 2 9
300 73.33 0.880 0.000 0.000 -0.120 22 0 3 5 300 66.67 0.864 0.095 0.172 -0.042 18 2 3 7
AMC23 0 97.50 - - - - - - - - MATH500 0 99.20 - - - - - - - -
30 97.50 1.000 0.000 0.000 0.000 39 0 0 1 30 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
60 97.50 1.000 0.000 0.000 0.000 39 0 0 1 60 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
90 97.50 1.000 0.000 0.000 0.000 39 0 0 1 90 99.20 0.998 0.002 0.004 0.000 495 1 1 3
120 97.50 1.000 0.000 0.000 0.000 39 0 0 1 120 99.00 0.996 0.002 0.004 -0.002 494 1 2 3
150 97.50 1.000 0.000 0.000 0.000 39 0 0 1 150 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
180 97.50 1.000 0.000 0.000 0.000 39 0 0 1 180 99.00 0.998 0.000 0.000 -0.002 495 0 1 4
210 97.50 1.000 0.000 0.000 0.000 39 0 0 1 210 98.80 0.994 0.002 0.004 -0.004 493 1 3 3
240 97.50 1.000 0.000 0.000 0.000 39 0 0 1 240 98.60 0.992 0.002 0.004 -0.006 492 1 4 3
270 97.50 1.000 0.000 0.000 0.000 39 0 0 1 270 98.40 0.992 0.000 0.000 -0.008 492 0 4 4
300 97.50 1.000 0.000 0.000 0.000 39 0 0 1 300 97.80 0.984 0.002 0.004 -0.014 488 1 8 3
Minerva 0 62.50 - - - - - - - - Olympiad 0 88.00 - - - - - - - -
30 62.13 0.935 0.059 0.111 -0.006 159 10 11 92 30 86.67 0.963 0.022 0.043 -0.015 572 13 22 68
60 59.93 0.929 0.031 0.059 -0.040 158 5 12 97 60 86.81 0.960 0.027 0.053 -0.013 570 16 24 65
90 61.76 0.924 0.065 0.122 -0.011 157 11 13 91 90 86.07 0.961 0.017 0.034 -0.022 571 10 23 71
120 62.87 0.953 0.053 0.100 0.006 162 9 8 93 120 86.52 0.961 0.022 0.044 -0.016 571 13 23 68
150 62.50 0.947 0.053 0.100 0.000 161 9 9 93 150 85.48 0.955 0.017 0.034 -0.028 567 10 27 71
180 59.56 0.906 0.049 0.094 -0.045 154 8 16 94 180 84.00 0.944 0.011 0.021 -0.045 561 6 33 75
210 60.29 0.924 0.043 0.082 -0.034 157 7 13 95 210 82.81 0.926 0.016 0.032 -0.058 550 9 44 72
240 61.40 0.941 0.042 0.080 -0.017 160 7 10 95 240 82.81 0.931 0.011 0.021 -0.058 553 6 41 75
270 59.19 0.906 0.043 0.083 -0.051 154 7 16 95 270 81.33 0.912 0.013 0.025 -0.075 542 7 52 74
300 58.46 0.912 0.025 0.049 -0.063 155 4 15 98 300 79.70 0.897 0.009 0.018 -0.093 533 5 61 76
Support dynamics during RLVR training. Tab. 4 summarizes how support-dynamics metrics evolve during RLVR training. Across benchmarks, SRR remains consistently high, showing that RLVR reliably preserves previously correct reasoning paths, while NSCR gradually decreases, indicating a steady contraction of the model’s solution support. On harder datasets, mild fluctuations in NDR and SDS reflect transient exploration that is subsequently pruned. Overall, these statistics reveal that RLVR reshapes rather than monotonically improves the reasoning distribution. It reinforces a narrow set of stable trajectories over time, which explains both the early-stage gains and the late-stage degradation observed on more diverse math tasks.
Overall takeaway: RLVR as precision enhancer, not capability expander. Across model scales (1.5B–14B) and domains (math, non-math, multimodal), RLVR consistently behaves as a support-bounded optimizer. With SRR near one but NDR near zero, and uniformly negative NSCR, RLVR enhances precision by concentrating mass on known high-reward solutions but rarely discovers new reasoning paths. This aligns with the Temporal Forgetting effect (Li et al., 2025). Breaking the invisible leash may thus require explicit exploration strategies that deliberately seed probability mass into underrepresented solution regions.
@@ 4.3 When Empirical Support Expansion Occurs
(1) RLVR Recomposes Subskills the Base Model Already Possesses.
@@
(2) RLVR Corrects Prompt-Format Misalignment and Extracts Latent Ability.
A second driver is the sensitivity to the prompt format. In several expanding tasks, especially dice and boxnet, the base model demonstrates partial competence but fails to follow the response format required by the reward function. RLVR reshapes the distribution to follow instructions more effectively, unlocking capabilities that were previously present but suppressed. Consistent with this interpretation, the perplexity gaps in Tab. 2 remain modest in expansion cases, showing that the “new” completions are stylistic or formatting variants of reasoning patterns already accessible to the base model, not fundamentally novel solutions beyond its support.
@@
Metric Model AIME24 AMC23 MATH500 Minerva Olympiad Avg.
@@
However, expansion remains sharply limited. Across the 1.5B–14B models in Tab. 1, NDR never exceeds $0.04$, and NSCR remains non-positive. Thus, RLVR’s gains arise from amplifying low-probability but existing solution fragments or format-correct variants, but never from discovering solutions truly absent from the base distribution.
5 Entropy Reduction and pass@k Trade-off
5.1 Experimental Setup
To study how RLVR reshapes the sampling distribution, we examine the base model and RLVR with a medium sampling budget $k=32$ on the math reasoning benchmarks. We quantify two entropy metrics:
5.2 Results: Precision Gains, Entropy Dynamics, and Trade-offs
@@
Tab. 5 shows that RLVR consistently improves avg@32 across benchmarks, raising average performance from 54.5% to 65.4% for ProRL and from 43.0% to 61.3% for DAPO (Yu et al., 2025). However, this increased precision comes at a cost: RLVR systematically reduces answer-level entropy, indicating a collapse onto fewer distinct solutions and empirically validating our theoretical prediction that reward optimization sharpens output distributions around known modes, thereby reducing effective support coverage. Notably, intrinsically harder tasks, like AIME or Minerva, still exhibit higher absolute answer-level entropy for both the base and RLVR models, suggesting that challenging problems inherently foster broader solution spaces that require exploration over more diverse completions.
@@
Our empirical analysis reveals a trade-off in RLVR: it improves precision by amplifying high-reward outputs, but simultaneously narrows the diversity of global solutions. This limitation is especially consequential in domains that admit multiple valid answers or benefit from creative reasoning, underscoring the need for explicit exploration mechanisms or diversity-promoting strategies to complement standard RLVR. Moreover, the observed divergence between token- and answer-level entropy highlights the need for a more nuanced interpretation of stochasticity in reward-optimized models, showing that precision gains often come at the expense of global diversity, and that maintaining controlled variability is critical for sustaining effective exploration.